# Can LLMs Maintain Fundamental Abilities under KV Cache Compression?

### v1

@@ 5 ShotKV
Based on our comprehensive analysis, we find that current unified KV cache compression methods exhibit significant performance degradation on Long-Context Generation tasks. In this section, we introduce ShotKV, a decoding-time compression method designed to mitigate this degradation. Our approach is motivated by two key insights: (1) Figure 2 demonstrates that n-shot example prompts receive substantial attention in reasoning benchmarks, and (2) Observation 4 reveals that chunk-level compression is particularly effective for long-context arithmetic reasoning tasks. Based on these findings, we hypothesize that each shot example represents a coherent chunk of information, leading us to design ShotKV to preserve shot examples intact during decoding time.
@@ 5.1 Implementation
Given a prompt with $n$ shots and generated tokens, we define:
@@
where $s_{i}$ represents the $i$-th shot example containing $k_{i}$ tokens, and $\alpha_{t,h}^{l}$ denotes the attention score for token $t$ in head $h$ at layer $l$. Once the prefill phase KV cache is compressed, it remains fixed throughout the generation process.
Given a prefill compression ratio $r_{p}$, we prioritize shots with higher scores while ensuring the total number of preserved tokens does not exceed the KV cache limit. Specifically, shots are ranked by their scores and selected in descending order until reaching the compression budget $r_{p}\times|KV_{\text{prefill}}|$. This shot-level selection strategy helps maintain the semantic coherence of important examples while adhering to memory constraints.
@@
Here, $KV^{C}_{\text{prefill}}$ represents the compressed prefill KV cache, and $S_{\text{preserved}}$ represents the optimal subset of shots to be preserved after compression. The first equation aims to maximize the total importance score of selected shots, where $\{s_{1},...,s_{n}\}$ represents all available shots and $\text{Score}_{\text{prefill}}(s_{i})$ is the importance score of shot $s_{i}$ computed using attention weights as defined earlier. The second equation enforces the memory constraint: the total number of tokens ($k_{i}$) in the selected shots must not exceed the allocated budget, which is determined by the prefill compression ratio $r_{p}$ multiplied by the original KV cache size.
@@
Given a decoding compression ratio $r_{d}$, we select tokens with the highest scores to preserve. The compressed decoding KV cache $KV^{C}_{\text{decoding}}$ retains only the top-$k$ tokens where $k=r_{d}\times|KV_{\text{decoding}}|$, effectively maintaining the most influential context tokens while reducing memory usage:
@@
Finally, we combine the compressed prefill and decoding KV caches to form the total compressed KV cache:
@@ 5.2 Empirical Results
In this section, we evaluate ShotKV under two scenarios: many-shot GSM8K with multiple KV cache compression methods, and LongGenBench-GSM8K with three unified compression methods that optimize the KV cache during generation.
@@
For LongGenBench-GSM8K evaluation, we employ three state-of-the-art unified compression methods as baselines: StreamingLLM (Xiao et al., 2024), H2O (Zhang et al., 2023), and PyramidInfer (Yang et al., 2024). We conduct experiments using LLaMA-3-8B-Instruct (Dubey et al., 2024) on the LongGenBench-GSM8K benchmark (Liu et al., 2024d), maintaining consistent parameters with Observation 6 ($K=35$, $T=20$). For many-shot GSM8K experiments, we follow the configuration detailed in Observation 4.
@@
From the  Table 3, we can see that ShotKV achieves the best performance on LongGenBench-GSM8K, maintaining high performance at low compression ratios. Specifically, at a compression ratio of 40%, ShotKV achieves 47.33% accuracy, surpassing the full kv cache baseline (46.00%) and showing substantial improvements over other methods (32.66%-39.50%). And Table 2 shows that ShotKV also achieves the best performance on many-shot GSM8K, maintaining high performance at low compression ratios. Even at aggressive compression ratios (25%-30%), ShotKV maintains relatively stable performance (26.83%-38.33%), while other methods experience more severe degradation (6.33%-16.67%). This superior performance can be attributed to two key design choices: (1) the preservation of complete shot examples during compression maintains the semantic coherence necessary for mathematical reasoning, and (2) the separation of prefill and decoding phase compression allows for more flexible and task-appropriate token retention strategies. These results suggest that our shot-aware compression strategy is particularly effective for long-context generation tasks that require maintaining complex reasoning chains, such as mathematical problem-solving.
6 Further Discussion
Our comprehensive analysis of KV cache compression reveals several important implications and limitations that warrant further discussion:
Table 3: KV cache compression methods’ performance on Many-shot GSM8K
Method 100% 40% 30% 20% 10%
FullKV 0.8235 - - - -
StreamingLLM - 0.8037 0.7835 0.7537 0.7432
H2O - 0.7832 0.7932 0.7428 0.5127
PyramidKV - 0.7834 0.7934 0.7832 0.7037
SnapKV - 0.7935 0.8038 0.7934 0.6827
ChunkKV - 0.7832 0.7932 0.7835 0.7932
ShotKV(Ours) - 0.8107 0.8082 0.8057 0.8037
Trade-off between Memory Efficiency and Task Performance
While KV cache compression methods effectively reduce memory usage, our findings highlight a non-uniform impact across different tasks. This suggests that deployment strategies should carefully consider task-specific requirements when selecting compression ratios, potentially implementing adaptive compression based on the task type.
Implications for Model Design
The higher sensitivity of instruct-tuned models to compression raises questions about the relationship between model training objectives and compression robustness. Future research might explore training techniques that explicitly optimize for compression resilience while maintaining instruction-following capabilities.
Limitations
@@ 7 Conclusion
This paper presents a systematic study of KV cache compression’s impact on LLMs’ core capabilities, revealing several key findings: (1) Performance degradation is highly task-dependent, with arithmetic reasoning tasks showing particular sensitivity to aggressive compression; (2) Instruct-tuned models demonstrate higher sensitivity to compression compared to their base counterparts; (3) Shorter prompts are more vulnerable to compression effects; (4) Chunk-level compression strategies show superior performance on complex long-context reasoning tasks; (5) Long-context generation tasks are more sensitive to compression than long-context reasoning tasks.
@@
Based on these insights, we proposed ShotKV, a novel compression method that separately handles prefill and decoding phases while preserving shot-level semantic coherence. Our method demonstrates superior performance on long-context arithmetic reasoning tasks and long-context generation tasks, maintaining high accuracy even at low compression ratios.
These findings have important implications for the deployment of LLMs in resource-constrained environments and suggest several promising directions for future research, including: (1) Development of task-adaptive compression strategies; (2) Investigation of compression-aware training methods; and (3) Extension of compression techniques to other model architectures and modalities.
Impact Statement
Appendix A Additional Related work
@@
Recent work has explored various strategies for sharing KV caches across transformer layers. Layer-Condensed KV Cache (LCKV) (Wu & Tu, 2024) computes KVs only for the top layer and pairs them with queries from all layers, while optionally retaining standard attention for a few top and bottom layers to mitigate performance degradation. Similarly, You Only Cache Once (YOCO) (Sun et al., 2024) computes KVs exclusively for the top layer but pairs them with queries from only the top half of layers, employing efficient attention in the bottom layers to maintain a constant cache size. In contrast, Cross-Layer Attention (CLA) (Brandon et al., 2024) divides layers into groups, pairing queries from all layers in each group with KVs from that group’s bottom layer. MiniCache (Liu et al., 2024b) introduces a novel method that merges layer-wise KV caches while enabling recovery during compute-in-place operations, optimizing KV cache size. These methods illustrate various trade-offs between computation, memory usage, and model performance when sharing KV caches across transformer layers.
@@
Recent advances in prompt compression have yielded innovative approaches to information density optimization in natural language processing. Research by Wingate et al. (2022) demonstrates how soft prompting techniques can achieve higher information density per token. Building upon this foundation, AutoCompressor (Chevalier et al., 2023) leverages soft prompts to both condense input sequences and expand model context windows. Parallel developments by Zhou et al. (2023) and Wang et al. (2023) showcase iterative summarization strategies using LLMs, establishing persistent memory mechanisms particularly beneficial for narrative construction and conversational systems. The progressive development of the LLMLingua framework (Jiang et al., 2023, 2024; Fei et al., 2024) has advanced prompt compression capabilities across extended context processing, logical reasoning, and retrieval-augmented generation. Notable contributions from Fei et al. (2024) demonstrate effective context management through automated segmentation and semantic condensation using pre-trained language models.
@@
General tasks refer to evaluating the overall performance of LLMs under mathematical inference, logic reasoning, and common knowledge GSM8K (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2020) are the representative tasks. The former focuses on the step-by-step reasoning ability of mathematical problem solving while the latter covers assessment of common sense and expertise in multiple areas. Besides, MATH (Hendrycks et al., 2021) spans various mathematical fields, ranging from elementary algebra to calculus, aiming to improve the mathematical problem-solving capabilities of LLMs. Meanwhile, MathQA (Amini et al., 2019) is a large-scale dataset comprising approximately 37,000 multiple-choice questions with precise annotations, designed to enhance the interpretability and performance of LLMs. In addition, BBH (Suzgun et al., 2022), a subset of BIG-Bench (Srivastava et al., 2022), focuses on challenging tasks. BBH includes multi-step reasoning problems, highlighting the importance of Chain-of-Thought prompting in LLMs. Similarly, CSQA (Talmor et al., 2019) is a task that combines knowledge graph-based multi-step reasoning with conversational capabilities. CSQA emphasizes inference and context understanding grounded in knowledge graphs. Normally, the general tasks apply automatic evaluation metrics (e.g. multi-choice accuracy) to ensure comparability and standardization. However, optimization strategies like KV cache compression may introduce challenges in executing the mentioned tasks. Filtering and dropping of contexts are involved in the compression strategy which may lead to an intermediate inference steps missing. In addition, in tasks such as MMLU that are highly dependent on knowledge coverage, compression may weaken the model’s ability to capture long context or rare domain knowledge (Yuan et al., 2024).
@@
Security tasks focus on assessing the robustness and protections of LLMs against harmful content, including truthfulness (Lin et al., 2021), toxicity (Hartvigsen et al., 2022), and bias (Liang et al., 2021). Recently, researchers noticed the weakness of LLMs in adversarial prompts (Zhu et al., 2023), especially in generating illegal or inappropriate content under jailbreak prompts. Shen et al. (2024) analyze the jailbreak prompts in real cases to reveal the failure of model security mechanism under complex malicious input. Meanwhile, Deng et al. (2023) demonstrates the multilingual jailbreak makes model security in low-resource languages easier to bypass, significantly increasing the probability that users of low-resource languages will generate insecure content. Similar to general tasks, KV optimization techniques can cause the model to ignore potential security threats when dealing with jailbreak prompts, thereby improving the success rate of adversarial prompts (Li et al., 2024a).
@@
Code generation tasks test the capacities of LLMs to generate code, which not only requires that the model can generate syntactic code based on natural language description but also has certain logical reasoning abilities. HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) are the commonly used benchmarks. They measure the functional correctness of the model by testing the results of the code’s execution.
@@
Recent developments in evaluating long-context models have produced a comprehensive ecosystem of benchmarks, focusing on both comprehension depth and retrieval efficiency. In the comprehension domain, $\infty$-Bench (Zhang et al., 2024) has established new standards by crafting evaluation scenarios exceeding 100,000 tokens, while LongBench (Bai et al., 2023, 2025) introduced multilingual assessment frameworks spanning document comprehension, text synthesis, and programming tasks. Further enriching this landscape, ZeroSCROLLS (Shaham et al., 2023) and L-Eval (An et al., 2023) have expanded evaluation criteria to encompass real-world applications, particularly in query-based content summarization. The emergence of many-shot learning as a distinct paradigm for extended context processing (Agarwal et al., 2024) has added another dimension to this field. Notable contributions from LongGenBench (Liu et al., 2024d) have advanced evaluation methodologies by combining extensive response generation requirements with efficient, cost-effective quality metrics.
The development of retrieval-focused benchmarks has taken a distinct approach, predominantly utilizing constructed datasets that enable precise experimental control, particularly in managing input sequence lengths. This methodology helps neutralize variations in model performance stemming from differences in training approaches. Substantial research efforts have yielded specialized synthetic frameworks for assessing retrieval capabilities (Kamradt, 2023; Mohtashami & Jaggi, 2023; Li et al., 2023; Liu et al., 2024c; Hsieh et al., 2024), while concurrent investigations have revealed the broader implications of extended context processing for enhanced reasoning capabilities (Tay et al., 2021).

### v2

@@ 4 ShotKV
Our comprehensive empirical investigation in Section 3.3 has systematically revealed critical vulnerabilities in current KV cache compression approaches when applied to a diverse range of fundamental LLM capabilities. Key findings indicate that:
  • Specific task categories, notably Arithmetic Reasoning (Observation 1) and Long-Context Generation (Observation 6), exhibit pronounced performance degradation under aggressive compression.
  • The integrity of prompt information is paramount; tasks that derive significant benefits from ICL and CoT (Observation 5) or rely heavily on n-shot prompts (as evidenced by the attention patterns in Figure 2) are particularly susceptible to information loss from compression.
  • Preserving semantic coherence is crucial, with chunk-level strategies showing promise in complex reasoning tasks (Observation 4), suggesting that compressing or discarding tokens without regard to these semantic units can be detrimental.
These observations collectively underscore the limitations of existing unified compression methods, which often fail to preserve nuanced structured information embedded in prompts, thereby leading to documented performance drops. This necessitates a more discerning compression strategy that is acutely aware of the semantic and structural importance of prompt components, especially for tasks demanding intricate reasoning and extensive generation.
To address these multifaceted challenges identified by our empirical study, we introduce ShotKV, a novel decoding-time compression method. ShotKV is specifically designed to mitigate the observed performance degradation by strategically managing KV cache during the prefill and decoding phases. Our approach is founded on the principle that n-shot examples in prompts are not merely token sequences, but constitute coherent semantic chunks vital for effective reasoning (a concept supported by Figure 2 and Observation 4). We therefore design ShotKV to preserve these shot examples intact during the prefill phase, complemented by a distinct strategy for the decoding phase, aiming for robust performance, particularly on the sensitive tasks highlighted in our analysis.
@@ 4.1 Implementation
Given a prompt with $n$ shots and tokens generated, we define:
@@
where $s_{i}$ represents the $i$-th shot example containing $k_{i}$ tokens. The term $\alpha_{t,h}^{l}$ denotes the attention weight assigned by the query vector (corresponding to the first token to be decoded immediately following the prompt) to the key vector of a token $t$ within shot $s_{i}$, in attention head $h$ at transformer layer $l$. Once the prefill phase KV cache is compressed based on these scores, it remains fixed throughout the generation process.
Given a prefill compression ratio $r_{p}$, we prioritize shots with higher scores while ensuring that the total number of preserved tokens does not exceed the KV cache limit. Specifically, the shots are ranked by their scores and selected in descending order until they reach the compression budget $r_{p}\times|KV_{\text{prefill}}|$. This shot-level selection strategy helps to maintain the semantic coherence of important examples while adhering to memory constraints.
@@
Here, $KV^{C}_{\text{prefill}}$ represents the compressed KV cache for prefilling and $S_{\text{preserved}}$ represents the optimal subset of shots that should be preserved after compression. The first equation aims to maximize the total importance score of the selected shots, where $\{s_{1},...,s_{n}\}$ represents all available shots and $\text{Score}_{\text{prefill}}(s_{i})$ is the importance score of the shot $s_{i}$ calculated using attention weights as defined earlier. The second equation enforces the memory constraint: the total number of tokens ($k_{i}$) in the selected shots must not exceed the allocated budget, which is determined by the prefill compression ratio $r_{p}$ multiplied by the original KV cache size.
@@
Here, for a previously generated token $t$, $\alpha_{t,h}^{l}$ is similarly defined as the attention weight assigned by the query vector of the current token being generated to the key vector of token $t$, within head $h$ at layer $l$. Thus, $\text{Score}_{\text{decoding}}(t)$ represents the total attention received by token $t$ from the current generation step.
Given a decoding compression ratio $r_{d}$, we select the tokens with the highest scores to preserve. The compressed decoding KV cache $KV^{C}_{\text{decoding}}$ retains only the top-$k$ tokens where $k=r_{d}\times|KV_{\text{decoding}}|$, effectively maintaining the most influential context tokens while reducing memory usage:
@@
Finally, we combine compressed prefill and decoding KV caches to form the total compressed KV cache:
@@ 4.2 Empirical Results
In this section, we evaluate ShotKV under two scenarios: many-shot Arithmetic Reasoning with multiple KV cache compression methods, and LG-GSM8K with three unified compression methods that optimize the KV cache during generation.
@@
For LG-GSM8K evaluation, we employ three state-of-the-art unified compression methods as baselines: StreamingLLM [10], H2O [21], and PyramidInfer [26]. We conduct experiments using LLaMA-3-8B-Instruct [39] on the LG-GSM8K benchmark [38], maintaining consistent parameters with Observation 6 ($K=35$, $T=20$). For many-shot Arithmetic Reasoning experiments, we follow the configuration detailed in Observation 4.
Table 4: KV cache compression methods’ performance on Many-shot Arithmetic Reasoning
Method 100% 40% 30% 20% 10%
FullKV 82.35 - - - -
StreamingLLM - 80.37 78.35 75.37 74.32
H2O - 78.32 79.32 74.28 51.27
PyramidKV - 78.34 79.34 78.32 70.37
SnapKV - 79.35 80.38 79.34 68.27
ChunkKV - 78.32 79.32 78.35 79.32
ShotKV(Ours) - 81.07 80.82 80.57 80.37
@@
From the  Table 4, we can see that ShotKV achieves the best performance on LG-GSM8K, maintaining high performance at low compression ratios. Specifically, at a compression ratio of 40%, ShotKV achieves 47.33% accuracy, surpassing the full kv cache baseline (46.00%) and showing substantial improvements over other methods (32.66%-39.50%). And Table 3 shows that ShotKV also achieves the best performance on many-shot Arithmetic Reasoning, maintaining high performance at low compression ratios. Even in aggressive compression ratios (25%-30%), ShotKV maintains relatively stable performance (26.83%-38.33%), while other methods experience more severe degradation (6.33%-16.67%). This superior performance can be attributed to two key design choices: (1) the preservation of complete shot examples during compression maintains the semantic coherence necessary for mathematical reasoning, and (2) the separation of prefill and decoding phase compression allows for more flexible and task-appropriate token retention strategies. These results suggest that our shot-aware compression strategy is particularly effective for long-context generation tasks that require maintaining complex reasoning chains, such as mathematical problem solving.
@@ 5 Conclusion
This paper presents KVFundaBench, a benchmark for systematically evaluating the effects of KV cache compression on various fundamental LLM capabilities. Our findings reveal that performance degradation is highly task dependent, with arithmetic reasoning and long-context generation being particularly sensitive (Task-Dependent Degradation and Long-Context Generation Sensitivity). We also highlight that compression sensitivity is influenced by a confluence of factors, including inherent model characteristics such as training dynamics (Model-Type Robustness), prompt-level attributes like length (Prompt Length Vulnerability), and the reliance on in-context examples (Prompt-Gain Sensitivity). Crucially, we demonstrate the importance of preserving the semantic integrity of prompt components, especially at a chunk or shot level, for complex reasoning and generation tasks where current methods often struggle and where chunk-based approaches show promise (Chunk-Level Superiority).
@@
Based on these insights, we introduced ShotKV, a novel compression framework that distinctively manages prefill and decoding phases while prioritizing shot-level semantic coherence to mitigate information loss in sensitive tasks. ShotKV demonstrates superior performance, notably on long-context arithmetic reasoning and generation tasks, maintaining high accuracy even at aggressive compression ratios. The results of KVFundaBench and the efficacy of ShotKV underscore the potential for more nuanced compression strategies and suggest promising future research avenues.
Appendix A Related Work
Key–value Cache Optimization Techniques
KV cache is the core component in LLM inference, which avoids repetitive computations by caching Key and Value vectors. However, the cost of caching KV increases exponentially with the expansion of the model size and the length of the context [45]. Some approaches have been published to alleviate the problem. For example, KV Compression designs efficient content selection strategies to filter and manage tokens [21, 27]. Some methods identify important tokens by focusing on high attention allocation [22], while others optimize token selection by combining attention scores with value vector norms to improve importance evaluation [46]. Techniques like PyramidInfer reduce critical contexts layer by layer based on the distribution of attention scores [26], and StreamingLLM preserves attention sinks to maintain stable performance in extended sequences [10]. Researchers reduce storage costs by merging similar context representations and solving input disturbances caused by compression [47]. For example, CaM [47] works by integrating the KV cache to be dropped into the retained cache in proportion to the attention weight. In addition, Yao et al. [48] proposes CacheBlend to achieve a selective KV recompute. Only partial KVs of crucial tokens are updated to reduce the delay in the prefill stage and increase the throughput. In addition, the dynamic budget allocation method is also used to optimize the KV cache, which adjusts the resource allocation in real time according to the importance of the context, providing a balance between performance and efficiency in multitask inference scenarios [24, 49].Wu et al. [50] proposes a prefill-decoding separation strategy to optimize the KV cache compression.
Evaluation of LLMs’ Fundamental Abilities
Accurately evaluating the fundamental capabilities of large language models is crucial to understand their true potential and limitations. The evaluation typically spans across several key dimensions: world knowledge tasks like MMLU [33],BBH [51] assess models’ grasp of diverse domains through multiple-choice questions; commonsense reasoning tasks such as CSQA [35] evaluate inference and context understanding abilities; arithmetic reasoning benchmarks like GSM8K [34] test mathematical problem-solving capabilities through step-by-step reasoning; code generation tasks including HumanEval [36] measure the ability to generate functionally correct code; and safety evaluations using benchmarks like JailBreakV [37] assess models’ robustness against harmful content generation. Additionally, long-context benchmarks such as LongBench [30, 31] and Need-In-A-Haystack (NIAH) [32] aiming to evaluate models’ long-context summarization and retrieval capabilities. Furthermore, LongGenBench [38] evaluates the models’ ability to process and generate responses for extended input sequences. And recently, in-context many-shot learning has been recognized as a long-context reasoning paradigm [44], which considers the number of shots as a critical factor in the performance of LLM. Although these tasks typically employ automatic evaluation metrics for standardization, KV cache compression may introduce unique challenges, particularly in tasks requiring complex reasoning chains or extensive knowledge retrieval.
@@ Appendix A Related Work
Recent work has explored various strategies for sharing KV caches across transformer layers. The Layer Condensed KV Cache (LCKV) [52] computes the KV only for the top layer and pairs them with queries from all layers, while optionally retaining standard attention for a few top and bottom layers to mitigate performance degradation. Similarly, You Only Cache Once (YOCO) [53] computes KVs exclusively for the top layer but pairs them with queries from only the top half of layers, employing efficient attention in the bottom layers to maintain a constant cache size. In contrast, Cross-Layer Attention (CLA) [54] divides layers into groups, pairing queries from all layers in each group with KVs from that group’s bottom layer. MiniCache [55] introduces a novel method that merges KV caches in layering while enabling recovery during compute-in-place operations, optimizing the size of the KV cache. These methods illustrate various trade-offs between computation, memory usage, and model performance when sharing KV caches across transformer layers.
@@
Recent advances in prompt compression have yielded innovative approaches to information density optimization in natural language processing. Research by Wingate et al. [56] demonstrates how soft prompting techniques can achieve higher information density per token. Building upon this foundation, AutoCompressor [57] leverages soft prompts to both condense input sequences and expand model context windows. Parallel developments by Zhou et al. [58] and Wang et al. [59] showcase iterative summarization strategies using LLMs, establishing persistent memory mechanisms particularly beneficial for narrative construction and conversational systems. The progressive development of the LLMLingua framework [60, 61, 62] has advanced prompt compression capabilities across extended context processing, logical reasoning, and retrieval-augmented generation. Notable contributions from Fei et al. [62] demonstrate effective context management through automated segmentation and semantic condensation using pre-trained language models.
@@
General tasks refer to evaluating the overall performance of LLMs under mathematical inference, logic reasoning, and common knowledge. GSM8K [34] and MMLU [33] are representative tasks. The former focuses on the step-by-step reasoning ability of mathematical problem solving, while the latter covers assessment of common sense and expertise in multiple areas. Besides, MATH [63] spans various mathematical fields, ranging from elementary algebra to calculus, aiming to improve the mathematical problem-solving capabilities of LLMs. Meanwhile, MathQA [64] is a large-scale dataset comprising approximately 37,000 multiple-choice questions with precise annotations, designed to enhance the interpretability and performance of LLMs. In addition, BBH [51], a subset of BIG-Bench [65], focuses on challenging tasks. BBH includes multi-step reasoning problems, highlighting the importance of Chain-of-Thought prompting in LLMs. Similarly, CSQA [35] is a task that combines knowledge graph-based multi-step reasoning with conversational capabilities. CSQA emphasizes inference and context understanding grounded in knowledge graphs. Normally, the general tasks apply automatic evaluation metrics (e.g. multi-choice accuracy) to ensure comparability and standardization. However, optimization strategies like KV cache compression may introduce challenges in executing the mentioned tasks. Filtering and dropping of contexts are involved in the compression strategy which may lead to an intermediate inference steps missing. In addition, in tasks such as MMLU that are highly dependent on knowledge coverage, compression may weaken the model’s ability to capture long context or rare domain knowledge [66].
@@
Security tasks focus on assessing the robustness and protections of LLMs against harmful content, including truthfulness [67], toxicity [68], and bias [69]. Recently, researchers noticed the weakness of LLMs in adversarial prompts [70], especially in generating illegal or inappropriate content under jailbreak prompts. Shen et al. [71] analyze the jailbreak prompts in real cases to reveal the failure of model security mechanism under complex malicious input. Meanwhile, Deng et al. [72] demonstrates the multilingual jailbreak makes model security in low-resource languages easier to bypass, significantly increasing the probability that users of low-resource languages will generate insecure content. Similar to general tasks, KV optimization techniques can cause the model to ignore potential security threats when dealing with jailbreak prompts, thereby improving the success rate of adversarial prompts [73].
@@
Code generation tasks test the capacities of LLMs to generate code, which not only requires that the model can generate syntactic code based on natural language description but also has certain logical reasoning abilities. HumanEval [36] and MBPP [74] are the commonly used benchmarks. They measure the functional correctness of the model by testing the results of the code’s execution.
@@
Recent developments in evaluating long-context models have produced a comprehensive ecosystem of benchmarks, focusing on both comprehension depth and retrieval efficiency. In the comprehension domain, $\infty$-Bench [75] has established new standards by crafting evaluation scenarios exceeding 100,000 tokens, while LongBench [30, 31] introduced multilingual assessment frameworks spanning document comprehension, text synthesis, and programming tasks. Further enriching this landscape, ZeroSCROLLS [76] and L-Eval [77] have expanded evaluation criteria to encompass real-world applications, particularly in query-based content summarization. The emergence of many-shot learning as a distinct paradigm for extended context processing [44] has added another dimension to this field. Notable contributions from LongGenBench [38] have advanced evaluation methodologies by combining extensive response generation requirements with efficient, cost-effective quality metrics.
The development of retrieval-focused benchmarks has taken a distinct approach, predominantly utilizing constructed datasets that enable precise experimental control, particularly in managing input sequence lengths. This methodology helps neutralize variations in model performance stemming from differences in training approaches. Substantial research efforts have yielded specialized synthetic frameworks for assessing retrieval capabilities [32, 78, 79, 80, 81], while concurrent investigations have revealed the broader implications of extended context processing for enhanced reasoning capabilities [82].